Given a binomial random variable with n​ = 100 and p​ = 0.3​, estimate the​ Pr[X greater than or equals ≥20]
Answer #1

n = 100

p = 0.3

µ = n * p = 100 * 0.3 = 30

sd = sqrt(n * p * (1 - p)) = sqrt(100 * 0.3 * 0.7) = 4.583

P(X > 20) P(X > 19.5)

                         =P(\frac{X-\mu}{\sigma}>\frac{19.5-\mu}{\sigma})

                         =P(Z>\frac{19.5-30}{4.583})

                         = P(Z > -2.29)

                         = 1 - P(Z < -2.29)

                         = 1 - 0.011

                         = 0.9890 (ans)